# A Coalition Game for On-demand Multi-modal 3D Automated Delivery System

### v1

@@ 4.2.1 Graph Attention Encoder
According to Figure 5, the method comprises graph attention encoding part moving to the cooperative decoding for handling both fleets. Before the encoder, the time and distance-based hop neighbourhood for edges are incorporated to find a spatial-temporal correlation for aerial and terrestrial network embedding.
Figure 5: Overview of the methodology
The incorporation of attention layers builds upon the recent work Zhang et al. (2023b); Kool et al. (2018). An adjacency matrices of temporal distribution are defined as $A_{i,j}=1$ if the time window of the two customers is close and if the late time window $l_{i}$ of customer $x_{i}$ is nearest to late time window $l_{j}$ of customer $x_{j}$; otherwise, $\boldsymbol{A}_{ij}=0$. The temporal and spatial adjacency neighbourhood for a node $x_{i}$ is defined by ${NB}^{T}_{i}$ and ${NB}^{S}_{i}$, which limits the neighbourhood in adjacency matrices with a time window by a threshold $\zeta$ and distance threshold by $\mu$, for UAVs and ADRs respectively. In addition, a parameter called density, $\rho$, is defined to determine the probability of how dense the environment is, namely, the existence of an obstacle along two connecting nodes. A broad spectrum of customers’ time and space correlation can be trained by tuning these thresholds and testing various scenarios, such as sparse and dense urban environments. First, the initial embedding for each node and fleet edges is computed in Equation 8 and 9 through a linear layer.
@@
$\mathbf{h}_{i}^{0}=\begin{cases}\operatorname{BN}\left(\boldsymbol{W}_{1}(\mathbf{x}_{i};\mathbf{x}_{i+N})+\boldsymbol{b}_{1}\right),&\text{ if }i\in\left\{1,\ldots,N\right\},\\ \operatorname{BN}\left(\boldsymbol{W}_{2}(\mathbf{x}_{i})+\boldsymbol{b}_{2}\right),&\text{ if }i\in\left\{0,N+1,\ldots,2N\right\}\end{cases}$ (8)
$\mathbf{e^{d}}_{ij}=\operatorname{BN}\left(\boldsymbol{W}_{3}E_{ij}^{d}+\boldsymbol{b}_{3}\right),\\ \mathbf{e^{r}}_{ij}=\operatorname{BN}\left(\boldsymbol{W}_{4}E_{ij}^{r}+\boldsymbol{b}_{4}\right),\text{ if }i,j\in{NB}^{T}_{i}$ (9)
@@
where $\boldsymbol{W}_{1}$, $\boldsymbol{W}_{2}$ , $\boldsymbol{W}_{3}$, $\boldsymbol{W}_{4}$, $\boldsymbol{b}_{1}$, $\boldsymbol{b}_{2}$, $\boldsymbol{b}_{3}$, and $\boldsymbol{b}_{4}$ represents the learnable parameters and $BN(.)$ represent batch normalization. The graph attention network can assign different importance to the customers within the neighbourhood based on the node role through the attention mechanism (Lei et al., 2022) by computing the pairwise attention weight $a^{ij}$ at $l$th layer as in Equations 10 and 11. In this setup, we have defined the attention weight in a time dimension manner, capturing the vehicle’s travelling ability as well. In other words, each connecting edge weight in the graph attention network is considered as the relative time that can passed by a certain vehicle within the time window between two nodes. Therefore, $\hat{e^{d}}_{ij}=\big{|}e_{i}-l_{j}-\frac{d_{ij}}{v_{d}}\big{|}$ for UAVs and $\hat{e^{r}}_{ij}=\big{|}e_{i}-l_{j}-\frac{d_{ij}}{v_{r}}\big{|}$ exist for ADRs. Where $d_{ij}$ is the shortest distance between node $i$ and $j$, $v_{d}$ and $v_{r}$ are the UAV and ADR maximum velocities, respectively. This configuration has been a trade-off between the case of time-window and distance proximity, leading to better convergence for training dual complex encoding schemes by incorporating the network and vehicle specifications so each node can gain a meaningful, distinguished, and absorbing heterogeneous embedding.
By doing so, each pickup node would be influenced by the corresponding delivery and potential close-time proximity of other pickup nodes.
$\alpha_{ij}^{\ell}|_{P}=\displaystyle\frac{\exp\left(\sigma\left(\boldsymbol{g_{1}}^{\ell^{T}}\left[\boldsymbol{W_{1}}^{\ell}\left(\boldsymbol{h}_{i}^{(\ell-1)}\left\|\boldsymbol{h}_{j}^{(\ell-1)}\right\|\hat{\boldsymbol{e}}_{ij}\right)\right]\right)\right)}{\sum_{z\in{NB}^{T}_{i}\cup{NB}^{S}_{i}}\exp\left(\sigma\left(\boldsymbol{g_{1}}^{\ell T}\left[\boldsymbol{W_{1}}^{\ell}\left(\boldsymbol{h}_{i}^{(\ell-1)}\left\|\boldsymbol{h}_{z}^{(\ell-1)}\right\|\hat{e}_{iz}\right)\right]\right)\right)}\text{ if }(i\in P,j\in P\cup D)$ (10)
$\alpha_{ij}^{\ell}|_{D}=\frac{\exp\left(\sigma\left(\boldsymbol{g_{2}}^{\ell^{T}}\left[\boldsymbol{W_{2}}^{\ell}\left(\boldsymbol{h}_{i}^{(\ell-1)}\left\|\boldsymbol{h}_{j}^{(\ell-1)}\right\|\hat{\boldsymbol{e}}_{ij}\right)\right]\right)\right)}{\sum_{z\in{NB}^{T}_{i}\cup{NB}^{S}_{i}}\exp\left(\sigma\left(\boldsymbol{g_{2}}^{\ell T}\left[\boldsymbol{W_{2}}^{\ell}\left(\boldsymbol{h}_{i}^{(\ell-1)}\left\|\boldsymbol{h}_{z}^{(\ell-1)}\right\|\hat{e}_{iz}\right)\right]\right)\right)}\text{ if }(i\in D,j\in P\cup D)$ (11)
@@
$\widehat{\boldsymbol{h}}_{i}^{(1)}=\mathrm{BN}\left(\boldsymbol{h}_{i}^{(0)}+\sum_{k=1}^{K}\boldsymbol{W}_{3}^{k}\boldsymbol{h}_{i}^{\prime}\right)$ (12)
$\boldsymbol{h}_{i}^{(1)}=\operatorname{BN}\left(\widehat{\boldsymbol{h}}_{i}^{(1)}+\varphi\left(\widehat{\boldsymbol{h}}_{i}^{(1)}\right)\right)$ (13)
@@
$\boldsymbol{h}_{i}^{l}=\sum_{j\in{NB}^{T}_{i}\cup{NB}^{S}_{i}\cup N}a_{ij}|_{P\cup D}\boldsymbol{W}_{l}^{V}\boldsymbol{h}_{j}^{(\mathbf{l-1})}+\sum_{j\in{NB}^{T}_{i}\cup{NB}^{S}_{i}\cup P}a_{ij}|_{P}\boldsymbol{W}_{l}^{V}\boldsymbol{h}_{j}^{(\mathbf{l-1})}+\sum_{j\in{NB}^{T}_{i}\cup{NB}^{S}_{i}\cup D}a_{ij}|_{D}\boldsymbol{W}_{l}^{V}\boldsymbol{h}_{j}^{(\mathbf{l-1})}$ (15)
As the output of the attention mechanism to get the final and average embedding shown in Equation 16, note that ADRs and UAVs do not share parameters through the encoding mechanism and are embedded separately with unique network configurations. Figure 6 illustrates a detailed design of the encoder.
$\boldsymbol{h}_{j}=\frac{1}{n}\sum_{i\in P\cup D}(\boldsymbol{h}_{i}^{l})_{j}$ (16)
@@ 4.2.2 Decoder
In the decoder, the context embedding of each vehicle will be aggregated by fleet states concatenated with node embedding to get agent embedding as Equation 17.
$\boldsymbol{x}_{k}^{(a)}=\boldsymbol{v}_{k,t}+\boldsymbol{W}_{5}\cdot\left[\overline{\boldsymbol{h}}_{(N)};\boldsymbol{v}_{1,t};\boldsymbol{v}_{2,t};\ldots;\boldsymbol{v}_{K,t}\right],\quad\forall k\in N^{d},N^{r}.$ (17)
@@
We define the query vector as the agent embedding $\boldsymbol{x}_{k}^{(a)}$, key vectors and value vectors as the customer embedding $\boldsymbol{h}_{i}$, and utilize the attention mechanism to compute the importance $\lambda_{k,i}$ of each customer $i$ to agent $k$ as in Equation 18.
$u_{k,i}=(\boldsymbol{W}_{6}\cdot\boldsymbol{x}_{k}^{(a)})^{T}\cdot(\boldsymbol{W}_{7}\cdot\boldsymbol{h}_{i}),\quad\forall i\in N,\quad\forall k\in N^{d},N^{r}$ (18)
@@
Furthermore, the decoding process encodes the joint information embedding to a query (Zhang et al., 2022). To do so, compare it with each customer’s key to acquire the attention coefficient, representing the compatibility between vehicle $k$ and customer $i$ at time $t$, according to Equation 20.
$\tilde{h}_{k,i}=(\boldsymbol{W}_{8}\cdot h_{v,k})^{T}\cdot(\boldsymbol{W}_{9}\cdot h_{i}),\quad\forall i\in N,\quad\forall k\in N^{d},N^{r}.$ (20)
To guarantee that each vehicle would not select the same node, a global mask is used to handle such situations and other operational and delivery constraints; the masking procedure is used for both fleets in the probability of the selecting node $i$, which is noted in Equation 21:
@@
with Clip parameter of C, and $\boldsymbol{W}_{5},\boldsymbol{W}_{6},\boldsymbol{W}_{7},\boldsymbol{W}_{8},\boldsymbol{W}_{9}$ are learnable weight parameters. As a result, a tour can be generated by selecting vehicle-node pairs at every step. The output will be given to the encoder for re-encode the states to account for removing the visited node and re-embedding the edges states to update to features of the network; thereby, the decoder can be influenced by a variation on network embedding at the decoding step to select the superior pair. The detailed design of the decoder is depicted in Figure 7. In the assignment step, three strategies were adopted to test if a priority-based model in the decoding stage would perform superiorly in contrast to joint learning. The latter is when a node is assigned to the highest probability in two modes, whereas the former is for either of the modes; there is a priority in the assignment, and the UAV-priority case is at less risk due to versatility over ADRs in normal delivery situations.

### v2

@@ 4.2.1 Graph Attention Encoder
The graph attention is utilized as an encoder which takes the node features and dual edge features and uses multi-head attention to encode node-edge interactions based on spatial-temporal dependencies across multiple layers. This problem is governed by heterogeneous temporal and spatial feature distribution, where there are non-linear correlations between the locations and the customer order arrival time (pickup time), where distance-based edge embedding into node-to-node attention cannot learn the complex coupling relationship among visiting locations in the problem. In this regard, a novel encoder using an adjacency mask mechanism based on a temporal and spatial graph, and feature-aware edge-enhancement to the node embedding is designed. Figure 6 illustrates a detailed design of the encoder.
It is noted that the spatial graph is the terrestrial graph with the shortest path distance as the edge weight. On the other hand, the temporal graph is featured by the time window value of each node rather than its location. This graph takes the aerial graph nodes; having said that, the edges are each node’s time separation. In other words, the edge weights are the time window difference between nodes. The adjacency masking mechanism is provided to guide the multi-head attention layer to account for proximity nodes rather than irrelevant ones. The proximity nodes are defined based on the edge weight values in the graph. It can be defined for each node as a set of neighbour nodes with which there is an edge connection. Accordingly, the spatial and temporal graph proximity are determined as $NB^{S}_{i}=\left\{G^{r}\mid\left\|x_{i}-x_{{j}}\right\|_{2}\leq\mu,(i,j)\in N\right\}$ and $NB^{T}_{i}=\left\{G^{d}\mid|e_{i}-l_{j}|\leq\zeta,(i,j)\in N\right\}$, respectively, where $\|\cdot\|_{2}$ denotes the L2 norm. This way, local adjacency information of either graph is extracted to give more importance to nodes with more similar features. In this regard, an adjacency matrix is defined as $A_{i,j}=1$ if there is a link between node $i$ and $j$, and zero otherwise. Therefore, the temporal and spatial connection for a node $i$ is defined by $\left\{A_{ij}|j\in{NB}^{T}_{i}\right\}$ and $\left\{A_{ij}|j\in{NB}^{S}_{i}\right\}$, which limits the neighbourhood in adjacency matrices with a time window by a threshold $\zeta$ and distance threshold by $\mu$. In addition, a parameter called density, $\rho$, is defined to determine the probability of how dense the environment is, particularly, the probability of the existence of an obstacle along two connecting nodes. This is the parameter that can either make two aerial and terrestrial networks (road networks) the very same for $\rho=1$ or make a fully directional connection network (like free airspace) for $\rho=0$. A broad spectrum of customers’ time and location correlation is extracted by varying thresholds in addition to different network embeddings for sparse and dense urban environments. First, the initial embedding for each node and fleet edges is computed in Equation 28 and 29 through a linear layer. Depots, pickups, deliveries are encoded separately into the initial hidden embedding, $\mathbf{h}_{i}^{0}$, based on their features $\mathbf{s}$.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{h}_{i}^{0}=\begin{cases}\operatorname{BN}\left(\boldsymbol{W}_{0}(\mathbf{s}_{i})+\boldsymbol{b}_{0}\right),&\text{ if }i\in D^{\prime},\\ \operatorname{BN}\left(\boldsymbol{W}_{1}(\mathbf{s}_{i};\mathbf{s}_{i+n})+\boldsymbol{b}_{1}\right),&\text{ if }i\in P,\\ \operatorname{BN}\left(\boldsymbol{W}_{2}(\mathbf{s}_{i})+\boldsymbol{b}_{2}\right),&\text{ if }i\in D\end{cases}}$ (28)
Where $\boldsymbol{W}_{0}$, $\boldsymbol{W}_{1}$, $\boldsymbol{W}_{2}$ , $\boldsymbol{W}_{3}$, $\boldsymbol{W}_{4}$, $\boldsymbol{b}_{0}$, $\boldsymbol{b}_{1}$, $\boldsymbol{b}_{2}$, $\boldsymbol{b}_{3}$, and $\boldsymbol{b}_{4}$ represents the learnable parameters and $BN(.)$ represent batch normalization. Similarly, two linear projections for edge features of UAV and ADR networks are used.
$\mathbf{e^{d}}_{ij}=\operatorname{BN}\left(\boldsymbol{W}_{3}\hat{E^{d}}_{ij}+\boldsymbol{b}_{3}\right),\\ \mathbf{e^{r}}_{ij}=\operatorname{BN}\left(\boldsymbol{W}_{4}\hat{E^{r}}_{ij}+\boldsymbol{b}_{4}\right)$ (29)
@@
In this setup, we have defined the attention weight in a time dimension manner, capturing the vehicle’s travelling feature as well. In other words, each connecting edge weight in the graph attention network is considered as the relative time that can be passed by a certain vehicle mode within the time window between two nodes. Therefore, $\hat{E^{d}}_{ij}=\big{|}e_{i}-l_{j}-\frac{d_{ij}}{v_{d}}\big{|}$ for UAVs, $d_{ij}\in E^{d}$ and $\hat{E^{r}}_{ij}=\big{|}e_{i}-l_{j}-\frac{d_{ij}}{v_{r}}\big{|}$ exist for ADRs, $d_{ij}\in E^{r}$. Where $d_{ij}$ is the shortest distance between node $i$ and $j$, $v_{d}$ and $v_{r}$ are the UAV and ADR maximum velocities, respectively. This configuration has been a trade-off between the case of time-window and distance proximity, by incorporating the network and vehicle specifications, leading to a distinguished and agent feature-aware embedding.
Afterwards, a multiple GAT Convolution layer is utilized to aggregate the node and edge information to obtain the final node embedding. To capture the precedence constraint for pickup and delivery nodes, a joint encoding of pickup and delivery sets separately is fused into the attention mechanism. That is, two additional attention subsets specifically for the pickup and delivery network are aggregated to the node embedding. According to (Lei et al., 2022), the attention score between the node and edge embedding for all sets of nodes and edges can be determined by Equation 30.
$\alpha_{ij}^{\ell}=\displaystyle\frac{\exp\left(\sigma\left(\boldsymbol{g}^{\ell^{T}}\left[\boldsymbol{W}^{\ell}\left(\boldsymbol{h}_{i}^{(\ell-1)}\left\|\boldsymbol{h}_{j}^{(\ell-1)}\right\|{\boldsymbol{e}}_{ij}\right)\right]\right)\right)}{\sum_{z}\exp\left(\sigma\left(\boldsymbol{g}^{\ell T}\left[\boldsymbol{W}^{\ell}\left(\boldsymbol{h}_{i}^{(\ell-1)}\left\|\boldsymbol{h}_{z}^{(\ell-1)}\right\|e_{iz}\right)\right]\right)\right)}$ (30)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha_{ij}^{\ell}=\begin{cases}\alpha_{ij}^{\ell}|_{A}^{UAV},&\text{ if }z={NB}^{T}_{i},\;\;e_{ij}=\mathbf{e^{d}}_{ij}\\ \alpha_{ij}^{\ell}|_{A}^{ADR},&\text{ if }z={NB}^{S}_{i},\;\;e_{ij}=\mathbf{e^{r}}_{ij}\end{cases}\;\;\;\;\forall(i,j)\in N,\;\boldsymbol{W^{l}}=\boldsymbol{W_{A}},\;\boldsymbol{g^{l}}=\boldsymbol{g_{A}}}$ (31)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha_{ij}^{\ell}=\begin{cases}\alpha_{ij}^{\ell}|_{P}^{UAV},&\text{ if }z={NB}^{T}_{i},\;\;e_{ij}=\mathbf{e^{d}}_{ij}\\ \alpha_{ij}^{\ell}|_{P}^{ADR},&\text{ if }z={NB}^{S}_{i},\;\;e_{ij}=\mathbf{e^{r}}_{ij}\end{cases}\;\;\;\;\forall(i,j)\in P,\;\boldsymbol{W^{l}}=\boldsymbol{W_{P}},\;\boldsymbol{g^{l}}=\boldsymbol{g_{P}}}$ (32)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha_{ij}^{\ell}=\begin{cases}\alpha_{ij}^{\ell}|_{D}^{UAV},&\text{ if }z={NB}^{T}_{i},\;\;e_{ij}=\mathbf{e^{d}}_{ij}\\ \alpha_{ij}^{\ell}|_{D}^{ADR},&\text{ if }z={NB}^{S}_{i},\;\;e_{ij}=\mathbf{e^{r}}_{ij}\end{cases}\;\;\;\;\forall(i,j)\in D,\;\boldsymbol{W^{l}}=\boldsymbol{W_{D}},\;\boldsymbol{g^{l}}=\boldsymbol{g_{D}}}$ (33)
where $\boldsymbol{g_{A}},\boldsymbol{g_{P}},\boldsymbol{g_{D}}$ and $\boldsymbol{W_{A}},\boldsymbol{W_{P}},\boldsymbol{W_{D}}$ are learnable weight vectors and matrices, for every node in the graph, pickup, and delivery set, respectively. Subsequently, each subset of attention is used to compute the weight value fused embedding of each node using a non-shareable weight parameter of $\boldsymbol{W}_{V}^{l}$, in Equation 34 and 35, in a multi-head heterogeneous attention network in each layer for each mode, respectively.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{h}_{i}^{l}|^{UAV}=\sum_{j\in N\cap NB^{T}}a^{l}_{ij}|^{UAV}_{A}\boldsymbol{W}_{V}^{l}\boldsymbol{h}_{j}^{(\mathbf{l-1})}+\sum_{j\in P\cap NB^{T}}a^{l}_{ij}|^{UAV}_{P}\boldsymbol{W}_{Vp}^{l}\boldsymbol{h}_{j}^{(\mathbf{l-1})}+\sum_{j\in D\cap NB^{T}}a^{l}_{ij}|^{UAV}_{D}\boldsymbol{W}_{Vd}^{l}\boldsymbol{h}_{j}^{(\mathbf{l-1})}}$ (34)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{h}_{i}^{l}|^{ADR}=\sum_{j\in N\cap NB^{S}}a^{l}_{ij}|^{ADR}_{A}\boldsymbol{W}_{V}^{l}\boldsymbol{h}_{j}^{(\mathbf{l-1})}+\sum_{j\in P\cap NB^{S}}a^{l}_{ij}|^{ADR}_{P}\boldsymbol{W}_{Vp}^{l}\boldsymbol{h}_{j}^{(\mathbf{l-1})}+\sum_{j\in D\cap NB^{S}}a^{l}_{ij}|^{ADR}_{D}\boldsymbol{W}_{Vd}^{l}\boldsymbol{h}_{j}^{(\mathbf{l-1})}}$ (35)
Then, we calculate the attention weights $K^{\prime}$ times in a multi-head mechanism. The final weight value vector is the summation over all heads, and each weight for each layer and $k$th head is $W_{3}^{k}h^{\prime}_{i,k}$. The final multi-head attention is shown, followed by a BN layer in Equation 36. Afterwards, a feed-forward layer with a residual connection and BN layer is employed and demonstrated in Equations 37 and 38, respectively.
$\widehat{\boldsymbol{h}}_{i}^{(l)}=\mathrm{BN}\left(\boldsymbol{h}_{i}^{(l-1)}+\sum_{k=1}^{K^{\prime}}\boldsymbol{W}_{3}^{k}\boldsymbol{h}_{i,k}^{\prime}\right)$ (36)
$\boldsymbol{\tilde{h}}_{i}^{(l)}=\operatorname{BN}\left(\widehat{\boldsymbol{h}}_{i}^{(l)}+\varphi\left(\widehat{\boldsymbol{h}}_{i}^{(1)}\right)\right)$ (37)
@@
As a result, the node embedding in one GAT layer can be figured by the output of Equation 37, and the next layer embedding can be found after the multi-head attention calculation in 36. Note that this process is done twice for each Equation of 34 and 35 for UAV and ADR node-edge encoding, respectively. Finally, after $L$ convolution layers, the output of the encoder is two distinct sets of embeddings, $\tilde{h_{i}}^{L}|^{UAV}$ and $\tilde{h_{i}}^{L}|^{ADR}$ for $i\in N$, which is used to get the final and average embedding shown in Equation 39. Note that ADRs and UAVs do not share parameters through the encoding mechanism and are embedded separately with unique network configurations.
$\boldsymbol{\bar{h}}_{j}|^{UAV,ADR}=\frac{1}{N}\sum_{i\in N}(\boldsymbol{\tilde{h}}_{i}^{L}|^{UAV,ADR})_{j}$ (39)
@@ 4.2.2 Decoder
After the final node embedding for each mode, the decoder performs node assignment for each mode iteratively until all nodes are served or no feasible nodes remain. In the decoder, first, the vehicle features are concatenated and projected through linear layers, followed by a batch normalization layer. These states include vehicle features, load, accumulated travel time and remaining battery, concatenated with the current node embedding. Equation 40 shows the UAV and ADR initial feature embedding into a higher hidden dimensional representation, $\mathbf{v^{d}}$ and $\mathbf{v^{r}}$, respectively.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{v^{d}}=\operatorname{BN}\left(\boldsymbol{W}_{d}[v^{d};\boldsymbol{\tilde{h}}|^{UAV}]+\boldsymbol{b}_{d}\right),\\ \mathbf{v^{r}}=\operatorname{BN}\left(\boldsymbol{W}_{r}[v^{r};\boldsymbol{\tilde{h}}|^{ADR}]+\boldsymbol{b}_{r}\right)}$ (40)
Where $\boldsymbol{W}_{d}$, $\boldsymbol{W}_{r}$, $\boldsymbol{b}_{d}$, and $\boldsymbol{b}_{r}$ represents the learnable parameters. The context embedding of each vehicle will be aggregated by dynamic states of current vehicle features with aggregated global graph embeddings, $\boldsymbol{h}_{(N)}$, to get the agent-vehicle-context embedding via linear transformations. Equation 41 and 42 combine per-vehicle features with global context embeddings for each mode, respectively, at each step $t$ of decoding.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{x}_{k}^{(a)}|^{UAV}=\mathbf{v^{k}_{t}}+\boldsymbol{W}_{5}\cdot\left[\overline{\boldsymbol{h}}_{(N)}|^{UAV};\boldsymbol{v}^{1}_{t};\boldsymbol{v}^{2}_{t};\ldots;\boldsymbol{v}^{N^{d}}_{t}\right],\quad\forall k\in N^{d}}$ (41)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{x}_{k}^{(a)}|^{ADR}=\mathbf{v^{k}_{t}}+\boldsymbol{W}_{6}\cdot\left[\overline{\boldsymbol{h}}_{(N)}|^{ADR};\boldsymbol{v}^{1}_{t};\boldsymbol{v}^{2}_{t};\ldots;\boldsymbol{v}^{N^{r}}_{t}\right],\quad\forall k\in N^{r}}$ (42)
@@
We define the query vector as the concatenated agent embedding for both mode into $\boldsymbol{x}_{k}^{(a)}$, key vectors as the customer embedding ${\tilde{h}}_{i}$, and value vectors to utilize the multi-head attention mechanism to compute the compatibility scores between vehicles and nodes of $u_{k,i}$ for each customer $i$ to agent $k$ as in Equations 43.
${q}_{k}$ $=\boldsymbol{W}_{7}\cdot\boldsymbol{x}_{k}^{(a)},\quad\forall k\in N^{d},N^{r}$ (43)
$\kappa_{i}$ $=\boldsymbol{W}_{8}\cdot{\tilde{h}}_{i},\quad\forall i\in N$
${V}_{i}$ $=\boldsymbol{W}_{9}\cdot{\tilde{h}}_{i},\quad\forall i\in N$
$u_{k,i}$ $={q}_{k}^{T}\cdot\kappa_{i},\quad\forall i\in N,\quad\forall k\in N^{d},N^{r}$
@@
Furthermore, the joint information embedding to a query of nodes in a single-attention mechanism is used, followed by (Zhang et al., 2022). It compares with each customer’s key to acquire the attention coefficient, representing the compatibility between vehicle $k$ and customer $i$ at time $t$, according to Equation 45.
$\tilde{h}_{k,i}=(\boldsymbol{W}_{10}\cdot h_{v,k})^{T}\cdot(\boldsymbol{W}_{11}\cdot\tilde{h}_{i}),\quad\forall i\in N,\quad\forall k\in N^{d},N^{r}.$ (45)
To guarantee that each vehicle would not select the same node, a global mask is used to handle such situations and other operational and delivery constraints; the masking procedure is used for both fleets in the probability of selecting node $i$, which is noted in Equation 46:
@@
with Clip parameter of C, and $\boldsymbol{W}_{5},\boldsymbol{W}_{6},\boldsymbol{W}_{7},\boldsymbol{W}_{8},\boldsymbol{W}_{9},\boldsymbol{W}_{10},\boldsymbol{W}_{11}$ are learnable weight parameters. As a result, a tour can be generated by selecting vehicle-node pairs at every step. The detailed design of the decoder is depicted in Figure 7. In the assignment step, two strategies were adopted to test if a priority-based model in the decoding stage would perform superiorly in contrast to joint learning. The latter is when a node is assigned to the highest probability in two modes, whereas the former is for either of the modes; there is a priority in the assignment, the UAV and ADR priority case.